OpenStudy (anonymous):

On a map, Elmwood Drive passes through R(4,-11) and S(0,-9) Taylor Road passes through J(6,-2) and K(4,-5). If they are straight lines, are the two streets perpendicular? Explain.

OpenStudy (azureilai):

To do this, you find the slope of each line. m = (y1-y2)/(x1-x2) so for Elmwood drive it is ((-0)-(-11))/(0-4) And for Taylor road ((-5)-(-2))/(4-6) You solve the two equations, and if they come out to be opposite reciprocals of each other, the lines are perpendicular.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

im confused when i figured out the the equations for Elmwood I got 2/-4 and for Taylor i got -3/-2 and i stuck after that please help :)

OpenStudy (azureilai):

(-9)-(-11) = (-9+11) [ I made a mistake writing zero earlier, that probably why it was wrong] -9+11 = 3 0-4=-4 First slope+ 3/-4 (-5)-(-2) = -5+2 = -3 4-6=-2 second slope: -3/-2 = 3/2 The roads are not perpendicular because their slopes are not opposite reciprocals of each other.

OpenStudy (anonymous):

isn't -9+11=2

OpenStudy (azureilai):

oops yes, sorry I was thinking -9+12. But even if it was 2, the slope becomes 2/-4 which is 1/-2, it is still not the opposite reciprocal of 3/2.

OpenStudy (anonymous):

thanks you are a big help

OpenStudy (azureilai):

no problem.

OpenStudy (anonymous):

determine whether the graphs of the following equations are parallel or perpendicular. Explain. 3x-9y=9, 3y=x+12 ,2x-6y=12